# Mechanistic Interpretability of LoRA-Adapted Language Models for Nuclear Reactor Safety Applications

### v1

@@ 1 INTRODUCTION
The deployment of software in safety-critical systems within a nuclear power plant must adhere to stringent quality assurance (QA) requirements, which is fundamentally different from its use in non-critical applications. In the United States (U.S.), regulations such as 10 CFR 50, Appendix B, mandate a documented design basis, comprehensive testing, and full traceability of a system’s logic and behavior [6]. This regulatory framework presents a significant challenge for advanced artificial intelligence (AI) and LLMs. The inherent black-box nature of these models makes it difficult, if not impossible, to verify the internal reasoning path an LLM took to arrive at a recommendation, conflicting with QA standards. This has largely limited AI deployment in nuclear facilities for non-safety roles. Therefore, a new paradigm focusing on internal transparency is required to bridge the gap between the capabilities of modern AI and the safety standards of the nuclear industry.
This paper addresses this gap by proposing a methodology based on mechanistic interpretability to enhance the trustworthiness of LLMs in nuclear applications. The goal is to increase the transparency of these opaque models, making their internal reasoning processes examinable and validatable, a prerequisite for qualification in systems important to safety, as outlined in key industrial standards such as the Institute of Electrical and Electronics Engineers (IEEE) Std 7-4.3.2 [7], and their associated regulatory guidelines, such as the U.S. National Regulatory Commission (NRC) Regulatory Guide 1.152 [8]. Recent initiatives like the NRC’s Technology-Inclusive Regulatory Framework [9] recognize the potential benefits of advanced digital systems while emphasizing the need for robust safety assurance methodologies. We demonstrate this approach by adapting a general-purpose LLM [10] to the specific domain of Boiling Water Reactors (BWRs) using documentation from authoritative bodies like the International Atomic Energy Agency (IAEA) and Organisation for Economic Co-operation and Development (OECD)/Nuclear Energy Agency (NEA) [11, 12].
@@
The integration of advanced digital systems into nuclear operations is an ongoing effort in the industry. Previous studies have addressed the modernization of instrumentation and control systems [18], human factors considerations for AI applications in nuclear plants [19, 20], and software quality assurance for nuclear computational tools [21]. These foundational studies established the importance of verification and validation (V&V) processes for safety-critical nuclear software systems, underscoring the need for interpretable AI methods in nuclear applications.
@@ 2.2 Mechanistic Interpretability for Nuclear-Grade Assurance
2.3 Language Generation Evaluation Metric
In our evaluation, we focused our quantitative analysis on the BLEU score. Given the safety-critical nature of the nuclear domain, the precision, fluency, and linguistic fidelity of the model’s output are considered paramount. An answer that is factually complete but poorly or ambiguously can lead to operational errors. Because BLEU directly measures these characteristics by evaluating n-gram precision, it was selected as the most appropriate metric for validating the model’s performance in this specific context.
@@ 3.1 Dataset Construction
We constructed a bespoke Question-Answering dataset to train and evaluate the model in a realistic nuclear engineering context. The source material for this dataset was a collection of authoritative technical documents covering BWR systems, safety procedures, and simulator specifications from various international bodies. Specifically, we utilized documentation from the OECD NEA on severe accident mitigation [12], an IAEA report on BWR simulators [11], and system descriptions from the U.S. NRC Technical Training Center [26].
The Question-Answering pairs were generated using a custom Python script that leveraged OpenAI’s gpt-4o-2024-08-06 application programming interface (API) [27]. This process involved segmenting the source texts into manageable chunks and prompting the model to create relevant questions and their corresponding answers based exclusively on the text provided. The resulting dataset was divided using stratified sampling to ensure a balanced representation across the source documents in both the training and evaluation sets.
@@ 3.2 Model Fine-Tuning
The base model for our study was Gemma-3-1b-it, a state-of-the-art, general-purpose LLM [10]. For domain adaptation, we fine-tuned this model on the training split of our custom Question-Answering dataset using the LoRA [15] technique, as previously described. The LoRA configuration was applied to a comprehensive set of target modules within the model’s attention and feed-forward layers: q proj, k proj, v proj, o proj, gate proj, up proj, and down proj. The complete set of hyperparameters used in this process is listed in Table 1. This process resulted in the specialized model, referred to as ”LoRA” in our experiments.
@@ 3.3.1 Identifying Key Neurons
We first measured the average activation of each neuron in the final Multilayer Perceptron (MLP) layer for both the base Gemma model and our fine-tuned LoRA model across the entire evaluation dataset. The change in average activation ($\Delta_{\text{act}}=\text{AvgAct}_{\text{LoRA}}-\text{AvgAct}_{\text{Base}}$) was calculated for every neuron. Neurons exhibiting the largest absolute change were designated as key neurons for further analysis. Our study focused on the top six neurons with the largest absolute change in activation, a group that comprised five amplified (positive change) and one suppressed (negative change) neuron.
@@ 4.1 LoRA Fine-Tuning Induces Neuron Specialization
Our initial analysis focused on identifying neurons affected by the LoRA fine-tuning process. Figure 1 illustrates the average change in activation of the 12 neurons that exhibited the largest absolute shifts between the base and the LoRA-tuned models. A distinct pattern emerged: one set of neurons was consistently amplified (e.g., Neurons #1086, #560, #1066, #176, and #890), showing a large positive change in activation, while another set was suppressed (e.g., Neuron #941), showing a negative change. The full activation distributions for these neurons are presented as boxplots in Figure 2, which show a consistent upward and downward shifts in the median activation values for the amplified and suppressed neurons, respectively.
@@ 5.1 Interpretation of Key Findings: From Independent Experts to Neural Circuits
Our results provide compelling evidence for a circuit-based hypothesis regarding the encoding of domain expertise in a fine-tuned LLM. The finding that silencing most key neurons individually was not statistically significant while silencing the group collectively was strongly refutes a simple model in which each specialized neuron acts as an independent expert. Instead, it suggests that domain knowledge is a distributed property of a neural circuit, and these neurons act in concert to produce expert-level, domain-specific outputs.
@@ 5.2 Functional Differentiation of Key Neurons
  • Neurons #1066 & #1086: Key Concept Neurons. As the most influential ”hub neurons”, their function appears to be encoding and retrieval of core, domain-specific technical concepts and entities. When neuron #1066 was silenced, the model’s answer to a question about Swedish SAMG updates lost specific institutional names like ”Swedish Defence Equipment (SDEk)” and defaulted to a generic, uninformative phrase like ”a new version of the SAMG system”. This indicates that these neurons are critical in providing the factual terminological backbone of an expert’s answer.
  • Neuron #941 (Suppressed): The Generalist Language Suppressor. This neuron’s activation was decreased during LoRA fine-tuning, suggesting that it performs a function useful for general language but detrimental to technical accuracy. The Base model often produced conversational, non-committal answers (”You’ve hit on a really important and complex question!…”) and hallucinated incorrect acronyms (”BWR (Bushel-Wall Reactor)”). LoRA learned to suppress neuron #941 to inhibit these undesirable generalist tendencies, allowing more specialized neurons to generate concise, factual responses.
  • Neuron #176: Procedural Logic Neuron. This neuron seems to specialize in encoding the logical relationship between different procedures or states. When asked about the transition from EOP to SAG, silencing #176 caused the model to produce a nonsensical answer about an ””S” sequence”. However, the functional LoRA model correctly identified the condition for the switch. This suggests that neuron #176 is crucial for representing the ”if-then” logic inherent in operational guidelines.
  • Neuron #560: Over-aggressive Generalization Neuron. Paradoxically, silencing this neuron improved the BLEU score. A review of its outputs suggests that it attempts to make answers more concise but sometimes does so by omitting key details. For example, when asked about maintaining water levels for jet pumps, the standard LoRA model stated it was for operating ”effectively and efficiently,” whereas the #560-silenced model more accurately stated it was for operating ”effectively and to prevent damage to the pump.” Silencing this neuron removed a flawed simplification, resulting in a more precise and higher-scoring answer.
@@ 5.3 Nuclear Safety Implications of Neuron Silencing
The quantitative performance degradation and functional roles inferred above have direct and significant implications for nuclear safety. The failure modes observed when silencing these key neurons mirror the types of errors that could lead to serious operational incidents. A qualitative analysis of the model outputs revealed three primary failure modes with significant safety consequences, as detailed in Table 2.
@@
As the table illustrates, the identified neural circuit encodes not only isolated facts, but also the structured, safety-oriented logic that is essential for nuclear operations. The degradation of this circuit translates directly to an increase in operational risk, highlighting the necessity of verifying the integrity of such circuits before deployment.
Furthermore, the phenomenon illustrated in Figure 3, where the average answer length dramatically decreased as the BLEU score substantially increased, demonstrates that the LoRA-adapted model learned to be more succinct and effective, prioritizing factual precision over the verbosity of the base model.
@@ 6 Conclusion
The core contribution of this study is a pathway towards achieving nuclear-grade AI assurance. The ability to identify, monitor, and causally test knowledge-bearing neural circuits offers a practical solution to the AI V&V challenge, which has limited the deployment of advanced AI in safety-critical nuclear applications. This approach has profound implications for the nuclear regulatory framework. For instance, by establishing baseline activation maps of key circuits during qualification, as specified in standards like IEEE Std 7-4.3.2 [7], operators and regulators can continuously monitor the AI’s internal reasoning and flag deviations and perform targeted re-validation after plant modifications. This aligns with human factors principles (NUREG-0711 [20]) by making AI recommendations more trustworthy and comprehensible.

### v2

@@ 1 INTRODUCTION
The deployment of software in safety-critical systems within a nuclear power plant must adhere to stringent quality assurance (QA) requirements, which is fundamentally different from its use in non-critical applications. In the United States (U.S.), regulations such as 10 CFR 50, Appendix B, mandate a documented design basis, comprehensive testing, and full traceability of a system’s logic and behavior [6]. This regulatory framework presents a significant challenge for advanced artificial intelligence (AI) and LLMs. The inherent black-box nature of these models makes it difficult, if not impossible, to verify the internal reasoning path an LLM took to arrive at a recommendation, conflicting with QA standards. This has largely limited AI deployment in nuclear facilities for non-safety roles. Therefore, a new paradigm focusing on internal transparency is required to bridge the gap between the capabilities of modern AI and the safety standards of the nuclear industry. Accordingly, any plant-facing AI must provide traceable, auditable links from external requirements and source documents to its internal computations and final outputs.
The integration of advanced digital systems into nuclear operations is an ongoing effort in the industry. Recent work ranges from the modernization of instrumentation and control systems [7] to the application of LLMs for transient identification [8] and simulator assistance [9]. However, these applications do not resolve the inherent black-box nature of the models. In parallel, foundational safety principles have remained a constant focus, with studies addressing human factors considerations for AI applications in nuclear plants [10, 11], software quality assurance for nuclear computational tools [12], and the need for robust fault diagnosis frameworks [13]. These studies established the importance of verification and validation (V&V) processes for safety-critical nuclear software systems, underscoring the need for interpretable AI methods in nuclear applications.
This paper addresses this gap by proposing a methodology based on mechanistic interpretability to enhance the trustworthiness of LLMs in nuclear applications. The goal is to increase the transparency of these opaque models, making their internal reasoning processes examinable and validatable, a prerequisite for qualification in systems important to safety, as outlined in key industrial standards such as the Institute of Electrical and Electronics Engineers (IEEE) Std 7-4.3.2 [14], and their associated regulatory guidelines, such as the U.S. Nuclear Regulatory Commission (NRC) Regulatory Guide 1.152 [15]. Recent initiatives like the NRC’s Technology-Inclusive Regulatory Framework [16] recognize the potential benefits of advanced digital systems while emphasizing the need for robust safety assurance methodologies. We demonstrate this approach by adapting a general-purpose LLM [17] to the specific domain of Boiling Water Reactors (BWRs) using documentation from authoritative bodies like the International Atomic Energy Agency (IAEA) and Organisation for Economic Co-operation and Development (OECD)/Nuclear Energy Agency (NEA) [18, 19].
@@ 3.1 Dataset Construction
We constructed a bespoke Question-Answering dataset to train and evaluate the model in a realistic nuclear engineering context. The source material for this dataset was a collection of authoritative technical documents covering BWR systems, safety procedures, and simulator specifications from various international bodies. Specifically, we utilized documentation from the OECD NEA on severe accident mitigation [19], an IAEA report on BWR simulators [18], and system descriptions from the U.S. NRC Technical Training Center [29], all chosen for their comprehensive and trusted coverage of the BWR domain.
The Question-Answering pairs were generated using a custom Python script that leveraged OpenAI’s gpt-4o-2024-08-06 application programming interface (API) [30]. To ensure high quality and traceability, this process involved several key steps: First, the source texts were segmented into manageable chunks (approx. 800 characters). Second, structured prompts instructed the model to create relevant questions and their corresponding answers based exclusively on the text provided, a constraint designed to prevent factual hallucination. Finally, the model returned the output in a structured JSON format, embedding metadata like the source file for each pair to ensure full traceability. This process yielded a total of 1,304 Question-Answering pairs: 506 pairs for the SAMG documentation, 698 pairs from the simulator reports, and 100 pairs from the systems descriptions.
The resulting dataset was divided using stratified sampling based on the source document to ensure a balanced representation. This approach was critical because each source document covers a distinct sub-domain of BWR knowledge. Stratified sampling ensures these topics are proportionally represented in both the training and evaluation sets, preventing topical bias and enabling a more robust assessment of the model’s overall domain expertise. The dataset was split into training ($80\%$) and evaluation ($20\%$) sets, maintaining the stratified distribution across all source documents.
@@ 3.2 Model Fine-Tuning
The base model for our study was Gemma-3-1b-it, a general-purpose LLM [17]. For domain adaptation, we fine-tuned this model on the training split of our custom Question-Answering dataset using the LoRA [22] technique. We selected LoRA among PEFT methods because it freezes the base weights and trains only low-rank adapter matrices, which helps preserve pre-trained knowledge and reduces catastrophic forgetting while enabling efficient adaptation. This locality is concentrated in a small number of adapter parameters, making activation shifts easier to trace and analyze, and allowing targeted causal tests. The LoRA configuration was applied to a comprehensive set of target modules within the model’s attention and feed-forward layers: q proj, k proj, v proj, o proj, gate proj, up proj, and down proj. The complete set of hyperparameters used in this process is listed in Table 1. This process resulted in the specialized model, referred to as ”LoRA” in our experiments.
@@ 3.3.1 Identifying Key Neurons
The first step in our analysis was to identify the specific neurons responsible for encoding newly acquired domain knowledge. We a priori focus on the final Multilayer Perceptron (MLP) layer, which is the last nonlinear feature-synthesis block that writes into the residual stream immediately before readout. This proximity to the output makes it the most plausible location to detect adaptation-induced features with a favorable signal-to-noise ratio while avoiding a broad multiple-comparison search. We measured the average activation of each neuron in the final Multilayer Perceptron (MLP) layer for both the base Gemma model and our fine-tuned LoRA model across the entire evaluation dataset, and computed the average activation ($\Delta_{\text{act}}=\text{AvgAct}_{\text{LoRA}}-\text{AvgAct}_{\text{Base}}$).
To select a compact core set for causal analysis, we applied an elbow-style criterion to the sorted magnitudes $|\Delta_{\text{act}}|$ within this layer. After ranking all neurons by $|\Delta_{\text{act}}|$, we observed a clear change in slope around the sixth item. Operationally, we retained the smallest $k$ for which the marginal gain fell below $20\%-25\%$ of the first step. We therefore focused on the top six neurons as the most substantially altered components, comprising five amplified ($\Delta_{\text{act}}>0$) and one suppressed ($\Delta_{\text{act}}<0$) unit to probe both amplification and suppression effects.
@@ 3.3.3 Statistical Analysis
3.4 Performance Evaluation Metric
To quantitatively assess the impact of our causal interventions, we selected the BLEU score as the primary metric. In the safety-critical context of nuclear operations, linguistic precision is not merely a matter of style but a direct component of safety. An answer that is factually correct but phrased ambiguously or imprecisely can lead to operator misinterpretation and subsequent errors. The BLEU score, by calculating n-gram precision against a reference answer, directly measures the textual fidelity and fluency of the model’s output. This makes it a suitable proxy for the clarity and reliability of the information provided, which is a paramount concern in this domain. We compute sentence-level BLEU-4 for each question in our evaluation set and report the mean of these scores. BLEU does not capture factual accuracy, so we interpret it as a clarity and fluency proxy that complements our mechanistic and qualitative analyses.
@@ 4.1 LoRA Fine-Tuning Induces Neuron Specialization
Our initial analysis focused on identifying neurons affected by the LoRA fine-tuning process. Figure 1 provides a summary, showing the average change in activation ($|\Delta_{\text{act}}|$) for the 12 most affected neurons. A distinct pattern emerged: one set of neurons was consistently amplified (e.g., Neurons #1086, #560, #1066, #176, and #890), showing a large positive change in activation, while another set was suppressed (e.g., Neuron #941), showing a negative change. For a more detailed statistical view, Figure 2 presents boxplots comparing the full activation distributions of these same neurons for the base and LoRA models. These plots visually summarize the median, interquartile range, and overall spread of activations for each neuron, providing the granular statistical information requested.
@@ 5.1 Interpretation of Key Findings: From Independent Experts to Neural Circuits
Our results provide evidence consistent with a circuit-based hypothesis regarding the encoding of domain expertise in a fine-tuned LLM. The finding that silencing most key neurons individually was not statistically significant while silencing the six-unit set was significant, which refutes a simple model in which each specialized neuron acts as an independent expert. Instead, it suggests that domain knowledge is a distributed property of a neural circuit, and these neurons act in concert to produce expert-level, domain-specific outputs. This circuit-based view is not just an academic distinction. It is fundamental to safety assurance in a nuclear context. Verifying the model’s reliability now means validating the integrity of distributed neural circuits responsible for safety-critical knowledge, a more complex but necessary task than simply checking individual components.
@@ 5.2 Functional Differentiation of Key Neurons
  • Neurons #1066 & #1086: Key Concept Neurons. As the most influential ”hub neurons”, their function appears to be encoding and retrieval of core, domain-specific technical concepts and entities. When neuron #1066 was silenced, the model’s answer to a question about Swedish SAMG updates lost specific institutional names like ”Swedish Defense Equipment (SDEk)” and defaulted to a generic, uninformative phrase like ”a new version of the SAMG system”. This indicates that these neurons are critical in providing the factual terminological backbone of an expert’s answer. In a control room, this failure mode, losing specific terminology, could lead to dangerous ambiguity, such as referring to ‘the pump’ instead of the ‘High-Pressure Coolant Injection system’ during a critical procedure.
  • Neuron #941 (Suppressed): The Generalist Language Suppressor. This neuron’s activation was decreased during LoRA fine-tuning, suggesting that it performs a function useful for general language but detrimental to technical accuracy. The Base model often produced conversational, non-committal answers ”You’ve hit on a really important and complex question!…” and hallucinated incorrect acronyms ”BWR (Bushel-Wall Reactor)”. LoRA learned to suppress neuron #941 to inhibit these undesirable generalist tendencies, allowing more specialized neurons to generate concise, factual responses. In practice, suppressing this neuron may reduce conversational filler and incorrect acronyms, thereby improving technical accuracy and conciseness.
  • Neuron #176: Procedural Logic Neuron. This neuron seems to specialize in encoding the logical relationship between different procedures or states. When asked about the transition from EOP to SAMG, silencing #176 caused the model to produce a nonsensical answer about an ””S” sequence”. However, the functional LoRA model correctly identified the condition for the switch. This suggests that neuron #176 is crucial for representing the ”if-then” logic inherent in operational guidelines. A malfunction in this neuron represents a direct operational threat, as it could cause the model to recommend an incorrect action sequence or fail to recognize the entry conditions for a critical procedure like the transition from EOP to SAMG.
  • Neuron #560: Over-aggressive Generalization Neuron. Paradoxically, silencing this neuron improved the BLEU score. A review of its outputs suggests that it attempts to make answers more concise but sometimes does so by omitting key details. For example, when asked about maintaining water levels for jet pumps, the standard LoRA model stated it was for operating ”effectively and efficiently,” whereas the #560-silenced model more accurately stated it was for operating ”effectively and to prevent damage to the pump.” Silencing this neuron removed a flawed simplification, resulting in a more precise and higher-scoring answer. This highlights a subtle but important safety concern: an AI that over-simplifies for the sake of conciseness could omit critical warnings or conditions, such as the risk of pump damage.
@@ 5.3 Nuclear Safety Implications of Neuron Silencing
The quantitative performance degradation and functional roles inferred above have direct and significant implications for nuclear safety. The failure modes observed when silencing these key neurons mirror the types of errors that could lead to serious operational incidents. A qualitative analysis of the model outputs revealed three primary failure modes with significant safety consequences.
  • (1) Emergency Response Procedure Failures. As illustrated by silencing neuron #176, disruption of procedural logic can produce incorrect EOP to SAMG transition recommendations, risking mistimed or mis-sequenced actions during accident mitigation.
  • (2) System Identification and Terminology Loss. Silencing neurons #1066 and #1086 led to loss of specific institutional names (e.g., “Swedish Defense Equipment”), and neuron #941 produced incorrect acronyms (e.g., “Bushel-Wall Reactor”), creating communication hazards in control room operations.
  • (3) Equipment Protection Logic Degradation. In the case of neuron #560, over-simplification omitted pump-damage-prevention conditions, posing subtle but important risks where essential cautions may be excluded.
    These failure patterns are detailed in Table 2.
@@
As Table 2 illustrates, the identified neural circuit encodes not only isolated facts, but also the structured, safety-oriented logic that is essential for nuclear operations. The degradation of this circuit may increase operational risk, highlighting the necessity of verifying the integrity of such circuits before deployment.
Furthermore, the phenomenon illustrated in Figure 3, where the average answer length decreased as the BLEU score increased, is consistent with improved clarity and terminological fidelity in the LoRA-adapted model relative to the base model. Although BLEU does not measure factual accuracy, the joint trend suggests more concise and precise language use. Two concrete, safety-relevant implementation scenarios are suggested by these findings. Decision support during plant transients: Recommendations could be accompanied by an interpretation dashboard indicating which key-neuron circuits are active for the current prompt, potentially improving operator trust and situational awareness. Operator training in a simulator: Circuit-activation signals could be used to compare a trainee’s response with the model’s reference pattern and to deliver targeted, context-aware feedback beyond simple correctness.
@@ 6 Conclusion
The core contribution of this study is a pathway towards achieving nuclear-grade AI assurance. The ability to identify, monitor, and causally test knowledge-bearing neural circuits offers a practical solution to the AI V&V challenge, which has limited the deployment of advanced AI in safety-critical nuclear applications. This approach has important implications for the nuclear regulatory framework. In our case study, analysis of individual neuron functions indicates that specific failure patterns can be traced to identifiable neural components: key concept encoding neurons (#1066, #1086), procedural logic representation neuron (#176), and protective-condition handling neuron (#560). This mechanistic understanding could inform targeted acceptance criteria, such as verifying terminology precision for concept neurons or testing procedural transition logic for sequence-critical neurons, thereby establishing candidate safety benchmarks for AI system qualification. For instance, establishing baseline activation maps of key circuits during qualification, consistent with standards like IEEE Std 7-4.3.2 [14], could enable operators and regulators to continuously monitor the AI’s internal reasoning and flag deviations and perform targeted re-validation after plant modifications. This aligns with human factors principles (NUREG-0711 [11]) by making AI recommendations more trustworthy and comprehensible.